OpenStudy (anonymous):

sum_{n=1}^{infty} [sqrt(n^4+1)/(n^3+n^2)] with steps

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \sqrt(n^4+1)/(n^3+n^2)\]

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \sqrt{\frac{n^4+1}{n^3+n^2}}\]

OpenStudy (anonymous):

no the square root sign is only for the numerator

OpenStudy (anonymous):

oh ok..

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac {\sqrt{n^4+1}}{n^3+n^2}\]

OpenStudy (anonymous):

yes that's it

OpenStudy (anonymous):

thanks, i know how do do that but i need the steps of how to do it